Find the angular magnification produced by a simple microscope of
focal length 5cm when used not in normal adjustment.
1
Expert's answer
2020-04-06T09:07:03-0400

As per the question,

Focal length (f)= 5cm

least distance for distinct vision (D)=25 cm

Angular magnification for simple microscope ="1+\dfrac{D}{f}"

"=1+\dfrac{25}{5}"

"=1+5"

"=6"